OpenStudy (anonymous):

In square ABCD, AE=3x+5 and BD=10x+2. What is the length of segment AC? answer choices are.... A.2 B.11 C.22 D.32

OpenStudy (anonymous):



OpenStudy (anonymous):

D

OpenStudy (anonymous):

Thank you !

OpenStudy (ranga):

In a square, the diagonals are the same length. Therefore, AC = BD BD = 10x + 2 (given). So AC = 10x + 2 Also AE is half of AC or two times AE = AC AE = 3x + 5 AC = 2 * AE 10x + 2 = 2(3x + 5) 10x+2 = 6x+10 10x - 6x = 10 - 2 4x = 8 x = 2 So AC = 10x+2 = 10(2) + 2 = 20 + 2 = 22. Answer: C.

OpenStudy (anonymous):

Yea hes right